OpenStudy (anonymous):

Solve the following equation. 3(y – 5) + 2 = 5 (1 point) y = 4 y = 7 y = –4 y = 6

OpenStudy (anonymous):

does anybody know this problem

OpenStudy (anonymous):

Here:

OpenStudy (anonymous):

Solve the following equation. 3(y – 5) + 2 = 5 (1 point) y = 4 y = 7 y = –4 y = 6 18. Solve the following equation. 75 = –5(–3 – 4x) (1 point) x = 3 x = –4 x = –3 x = 5 19. Solve the following equation. –2x + 4 = 2(4x – 3) – 3(–8 + 4x) (1 point) x = 7 x = 2 x = –7 x = 3

OpenStudy (anonymous):

does anyone know these

OpenStudy (anonymous):

18-

OpenStudy (anonymous):

19-

OpenStudy (anonymous):

Are you not seeing the photos? Lol

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve the following equation. u + 2 = 4

OpenStudy (anonymous):

is the answer 2 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Solve the following equation. –5 = s – 3

OpenStudy (anonymous):

Add the 3 to both sides

OpenStudy (anonymous):

So it will be: -2=s

OpenStudy (anonymous):

Or you could flip it around if you are used to having the variable on the right side : s=-2 It's the same thing

OpenStudy (anonymous):

*sorry the variable on the left side

OpenStudy (anonymous):

Solve the following equation. five over thirteen t = –9

OpenStudy (anonymous):

Is the t only with the 13, or with the whole fraction?

OpenStudy (anonymous):

A or b

OpenStudy (anonymous):

a

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Solve the following equation. 13 + w over 7 = –18